OpenStudy (anonymous):

Factor completely: x^2 - 36

OpenStudy (anonymous):

\[x^2 - 36 = x^2 - 6^2 = (X+6)(X-6)\]

OpenStudy (anonymous):

Seeing the problem, means that you will have this type of problem\[(a+b)(a-b)\]